Convert the given fractional numbers to percent.
 3/40

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

In order to convert a fraction into a percentage multiply the fraction by 100 and put the percent sign %.
= (3/40) × 100 %
= 300/40 %
= 30/4 %
= 7.5%

​Thanks